OpenStudy (anonymous):

Use the graph of y = e^x to evaluate e^1.7 to four decimal places. @whpalmer4 and there is no graph given A. 5.4739 B. 4.6211 C. 2.7183 D. 0.1827

OpenStudy (anonymous):

the graph is y=e^x ?

OpenStudy (anonymous):

yea silly me ..

OpenStudy (whpalmer4):

No problem, I'll make you one :-)

OpenStudy (anonymous):

ok so how do I solve this problem ?

OpenStudy (whpalmer4):

Okay, x-axis is the x value, I've put in vertical lines at x = 1.5, 1.6, 1.7, 1.8, go up at x=1.7 to the curve, then go left to the y-axis and figure out which value of your four choices is the spot where the y-axis is intercepted by the horizontal line...

OpenStudy (whpalmer4):

you aren't really estimating it to 4 places, you're picking from a set of estimates the closest one, and they just happen to provide 4 places, which is kind of bogus, but whatever...

OpenStudy (wolf1728):

You can easily use a calculator to evaluate 'e' or 2.718281828 to the power of 1.7 and it is 5.4739473902 Looks as if the answer is A

OpenStudy (whpalmer4):

Here's a slightly different graph: here I've marked the answer choices as horizontal lines, and you can see which one intersects the curve at x = 1.7...

OpenStudy (anonymous):

Okay I guess I understand.. Nope I don't :(

OpenStudy (whpalmer4):

it's just reading values off a graph....you want to know the value of the function \[f(x) = e^x\] at \(x=1.7\) so you draw a perpendicular line across the x-axis at \(x=1.7\) until it intersects the curve. At that point, you draw a perpendicular line over to the y-axis, and read the value off the scale. Now, probably you are confused as to how the heck you are supposed to get 4 digits reading off a scale that is labeled 1 2 3 4 5 6 7! You just have to see that "oh, it crosses somewhere between 5 and 6, so the answer must be the one that is between 5 and 6, which is A) 5.4739"

OpenStudy (anonymous):

Ok understood, clarification

OpenStudy (whpalmer4):

If they weren't giving you values to choose from, you couldn't possibly answer and give 4 digits — you could estimate that it is about 5.5 or so, and that's it.

OpenStudy (anonymous):

Great thanks for all the help hope you can help me again some time, Goodnight ^_^

OpenStudy (whpalmer4):

Bad problem-writing, if you ask me! Either confuses the student, or gives the mistaken impression that you could accurately read the graph to 4 digits... You're very welcome, and I hope to help you again :-) Good night!

OpenStudy (wolf1728):

Hello Shay

OpenStudy (anonymous):

Hey @wolf1728 :)

OpenStudy (wolf1728):

Hello - still need help?

OpenStudy (anonymous):

Yup you will be tagged in 2 seconds